OpenStudy (anonymous):

Would this be the formula for a square? Area of A = a2 = 20m × 20m = 400m2

OpenStudy (anonymous):

the formula of a square would just be: Area of A = a^2

Directrix (directrix):

A sqare of side a = 20 m would have area 400 square mm so if that is what you are doing, then, yes, you are correct.

OpenStudy (anonymous):

It ask me to explain each variable in the formulas of a square, rectangle, and triangle.

Directrix (directrix):

In the square formula, A = a^2, the variable a represents the side of the square. A square has all sides the same length. Bring the other two formulas over here and we'll do them.

OpenStudy (anonymous):

This is for a triangle: Area of B = ½b × h = ½ × 20m × 14m = 140m2

Directrix (directrix):

Triangle: Area = 1/2 b x h where b is a base of the triangle and h represents the length of the altitude drawn to that base. If b = 20 m and h = 14 m , the area of the triangle is 1/2 (20) (14) = 140 m^2. Leave this "B" off and write "Area of triangle" = Now, get the rectangle formula and bring it here.

OpenStudy (anonymous):

The formula for perimeter of a rectangle is: P = 2L + 2W.

Directrix (directrix):

Was this supposed to be area of a rectangle? I may be mixed up. ------- P = 2L + 2W is the formula for the perimeter of a rectangle. L represents the length of the rectangle and W represents the width. Some people write the formula as A = L + W + L + W. That equals 2L + 2W for the perimeter of the rectangle.

OpenStudy (anonymous):

yes

Directrix (directrix):

Then, get the area of a rectangle formula, bring it over here, and we'll crank it out. Perimeter is the distance around the rectangular region but area is how much region (area) in included inside the rectangle.

OpenStudy (anonymous):

ok

Directrix (directrix):

After that, then this question. I think something is missing. A water fountain with a 4 ft. by 4 ft square base is placed on a square grassy park area that is x ft on a side. To determine the amount of grass seed needed for the lawn, find a polynomial for the grassy area.

OpenStudy (anonymous):

Would this be it Area=10.0x5.0=50.0

Directrix (directrix):

It might be an example but where is the formula for the area of a rectangle.A = ?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

Area = w × h w = width h = height

Directrix (directrix):

 Fountain base area is A = s^2 = 4^2 = 16 Grass Plot area is A = s^2 = x^2 Region in park but not covered by fountain" Y = x^2 - 16